5/6 divided by 1/2 x 1/5
Answer by checkley77(12569)

5/6 divided by 1/2 x 1/5
(5/6)/(1/2*1/5)
(5/6)/(1/10)
5/6*10/1
50/6=25/3=8 1/3=8.33333 ans.